Circular No. 7472

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS Further to IAUC 7459, D. Hammer provides measurements of six

comets found in archival images obtained in 1997 with the SOHO C2 telescope only. M. Oates found C/1997 J3, C/1997 J4, and C/1997 K3; X. Leprette found C/1997 L5 and C/1997 M3; and Hammer found C/1997 K4. The discovery observations for these Kreutz sungrazers are given below, while MPEC 2000-P18 contains the fully reduced observations and orbital elements (by B. G. Marsden).

1997 UT  α_{2000}  δ_{2000}  Comet
May 10.397 3^{h}08.^{m}7 +16^{o}16′ C/1997 J3
10.397 3 11.8 +16 28  C/1997 J4
25.905 4 09.5 +19 34  C/1997 K3
31.777 4 34.2 +20 20  C/1997 K4
June 12.097 5 16.7 +21 34  C/1997 L5
28.603 6 24.5 +21 58  C/1997 M3

S/1999 J 1 B. Gladman, Observatoire de la Côte d’Azur; and H. Boehnhardt,

European Southern Observatory, report photometry of this satellite, taken with the 8.2-m ANTU Very Large Telescope on July 29: B = 21.43 ± 0.09, V = 20.73 ± 0.04, R = 20.21 ± 0.03, I = 19.87 ± 0.06.

COMET C/1999 S4 (LINEAR) M. Kidger, Instituto de Astrofísica de Canarias, reports analysis of

continued observations of C/1999 S4 at La Palma: “A 100-s exposure in R, taken by R. Corradi and N. O’Mahoney (Isaac Newton Group of Telescopes) on Aug. 1.9 UT with the Wide Field Camera on the 2.5-m Isaac Newton Telescope, shows no evidence of a nuclear condensation or subnuclei within the coma. The seeing was 1′′.5. It is thus highly improbable that any fragments of the nucleus of significant size exist within the coma. The coma shows a structure similar to that reported previously (IAUC 7467, 7468), with a well-defined sunward boundary to the coma similar to the point of a lance, although the brightest part of the coma is now displaced > 1′ tailwards. The tail can be traced at least 20′.”

2000 August 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green